### v1

@@ 9 Tables and Figures
$\hat{\theta}_{1}$ $\hat{\theta}_{2}$ $\hat{p}_{1}$ $\hat{p}_{fb,d}$ $\hat{p}_{sb,d}$ $\hat{\theta}_{1}$ $\hat{\theta}_{2}$
Estimates 0.0218 0.0018 0.3663 0.1124 0.3217 0.0115 0.0016
Standard errors 0.0017 0.005 0.0145 0.0119 0.0136
Number of observations 11,270 942
Table 7: Data and Model Moments: Targeted Moments
Data Moments Model moments Difference
$\bar{Y}_{d}^{d}-\bar{Y}_{d}^{b}$ 1.7028 1.7169 -0.0141
Non-educated parents $\bar{Y}_{sb}^{d}-\bar{Y}_{fb}^{d}$ 0.2754 0.4502 -0.1748
$\bar{Y}_{sb}^{s}-\bar{Y}_{fb}^{s}$ 0.3682 0.1616 0.2066
Proportion of educated firstborn 0.3642 0.3678 0.0036
Proportion of educated firstborn daughters 0.1116 0.1156 0.004
Proportion of educated second born daughters 0.3177 0.3177 0
$\bar{Y}_{d}^{d}-\bar{Y}_{d}^{b}$ 1.56 1.55 0.01
College educated parents $\bar{Y}_{sb}^{d}-\bar{Y}_{fb}^{d}$ 2.61 2.76 -0.15
$\bar{Y}_{sb}^{s}-\bar{Y}_{fb}^{s}$ 2.47 2.88 -0.41
Proportion of educated firstborn
Proportion of educated firstborn daughters
Proportion of educated second born daughters
Table 8: Data and Model Moments: Non targeted Moments
Data Moments Model moments Difference
Average inequality 5.87 5.97 -0.1
Average inequality (only daughters households) 5.29 5.46 -0.17
Average inequality (only sons households) 5.71 5.87 -0.16
Average inequality (both gender households) 6.29 6.29 0
Average inequality ($\bar{q}<6$) 5.05 5.04 0.01
Non-educated parents Average inequality ($106\leq\bar{q}<$ 10) 8.46 7.56 0.9
Average inequality ($1310\leq\bar{q}<$ 13) 5.27 5.34 -0.07
Average inequality ($\bar{q}\geq 13$) 2.1 5.62 -3.52
Average inequality (intensive margin) 3.59 3.77 -0.18
Standard deviation of firstborn’s education 5.55 5.5
Standard deviation of second born education 5.19 5.29
Average inequality
Average inequality (only daughters households)
Average inequality (only sons households)
Average inequality (both gender households)
Average inequality ($\bar{q}<6$)
College educated parents Average inequality ($106\leq\bar{q}<$ 10)
Average inequality ($1310\leq\bar{q}<$ 13)
Average inequality ($\bar{q}\geq 13$)
Average inequality (intensive margin)
Standard deviation of firstborn’s education
Standard deviation of second born education
Table 9: Difference in Average education of men and women
Observed Unobserved
Average years of education
Male 5.35 4.66
Female 4.73 2.03
Difference 0.62 2.63
Proportion of married
Male 0.37 0.84
Female 0.38 0.88
Difference 0.01 -0.04
Quantity-quality trade-off in Benin
Table 10: Regression of education attainment dummy on number of children and other covariates
1998-2013 1970-1998 1960-1970
OLS IV OLS IV OLS IV
(Intercept) $0.74^{*}$ $0.71^{***}$ $0.44^{*}$ $0.46^{***}$ $0.16^{*}$ $0.62^{*}$
$[0.74;0.75]$ $(0.01)$ $[0.43;0.44]$ $(0.02)$ $[0.12;0.19]$ $(0.26)$
Family size:Urban $-0.01^{*}$ $-0.01^{***}$ $-0.00^{*}$ $-0.00$ $0.00$ $0.01$
$[-0.01;-0.01]$ $(0.00)$ $[-0.00;-0.00]$ $(0.01)$ $[-0.01;0.01]$ $(0.06)$
Family size:Rural $-0.02^{*}$ $-0.02^{***}$ $-0.00^{*}$ $-0.01^{*}$ $0.00$ $-0.16$
$[-0.03;-0.02]$ $(0.00)$ $[-0.00;-0.00]$ $(0.00)$ $[-0.00;0.01]$ $(0.10)$
Education of hh $0.02^{*}$ $0.02^{***}$ $0.03^{*}$ $0.03^{***}$ $0.05^{*}$ $0.05^{***}$
$[0.02;0.02]$ $(0.00)$ $[0.03;0.03]$ $(0.00)$ $[0.05;0.06]$ $(0.00)$
Female hh $0.09^{*}$ $0.10^{***}$ $0.11^{*}$ $0.11^{***}$ $0.15^{*}$ $0.03$
$[0.09;0.09]$ $(0.00)$ $[0.11;0.12]$ $(0.01)$ $[0.13;0.17]$ $(0.09)$
HWI $0.04^{*}$ $0.04^{***}$ $0.06^{*}$ $0.06^{***}$ $0.07^{*}$ $0.07^{***}$
$[0.04;0.04]$ $(0.00)$ $[0.06;0.06]$ $(0.00)$ $[0.06;0.08]$ $(0.01)$
Urban $0.02^{*}$ $0.04^{***}$ $0.13^{*}$ $0.11^{***}$ $0.16^{*}$ $-0.36$
$[0.02;0.02]$ $(0.01)$ $[0.12;0.13]$ $(0.03)$ $[0.12;0.19]$ $(0.30)$
Birth order $0.02^{*}$ $0.01^{***}$ $0.01^{*}$ $0.02^{***}$ $0.01$ $0.12$
$[0.01;0.02]$ $(0.00)$ $[0.01;0.01]$ $(0.00)$ $[-0.01;0.03]$ $(0.08)$
Female $-0.04^{*}$ $-0.04^{***}$ $-0.08^{*}$ $-0.08^{***}$ $-0.12^{*}$ $-0.14^{***}$
$[-0.04;-0.04]$ $(0.00)$ $[-0.09;-0.08]$ $(0.00)$ $[-0.14;-0.10]$ $(0.02)$
R2 $0.14$ $0.14$ $0.21$ $0.21$ $0.27$ $-0.05$
Adj. R2 $0.14$ $0.14$ $0.21$ $0.21$ $0.27$ $-0.05$
Num. obs. $2069182$ $2069182$ $403850$ $403850$ $6461$ $6461$
RMSE $0.40$ $0.43$ $0.40$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$ (or Null hypothesis value outside the confidence interval).
10 Appendix A: Additional Figures

### v3

@@ 5.1 Estimation Results
$\hat{\theta}_{1}^{\text{low}}$ $\hat{\alpha}_{1}^{\text{low}}$ $\hat{\theta}_{1}^{\text{ high}}$ $\hat{\alpha}_{1}^{\text{high}}$ $\hat{\theta}_{1}^{\text{low}}$ $\hat{\alpha}_{1}^{\text{low}}$
Estimates 0.0238∗∗ 0.0014 0.13∗∗ 0.017∗∗ 0.00759∗∗ 0.00045
Standard errors 0.0017 0.0004 0.0019 0.001 0.0006 0.0012
Number of observations 6187 5083 942
**∗ ∗ significant at 5% level of significance.
Result 1: For parents without formal education and with high financial, social and cultural constraints, parents’ perceived average utility at high school level of education is $\approx$ 11.5% higher for a second born child compared to firstborn child of the same gender.
The estimate of the marginal educational cost difference between firstborn and second born children suggests that, on average, for parents without formal education and with high financial, social and cultural constraints, the likelihood of the firstborn child being educated compared to a second-born child of the same gender is approximately 0.3663, which correspond to an average cost difference of $0.02$. Holding everything else equal, this cost difference translate into a utility gap of $\approx$ 11.5% for high school level of education.
Result 2: For parents without formal education and with high financial, social and cultural constraints, perceived average utility at high school level of education is $\approx$ 40% higher for the second born son compared to the firstborn daughter.
Among parents without formal education and high financial, social and cultural constraints, the marginal utility from an additional year of education is 33% high if given to a son compared to given to a daughter. After factoring in birth order, these parents’ perceived average utility at high school level of education is $\approx$ 40% higher for the second born son compared to the firstborn daughter. Their perceived average utility of graduating high school is $\approx$ 18% higher for the firstborn son compared to the second born daughter . Note that estimates for the extensive margin parameters ($\hat{\theta}_{1}^{\text{high}}$ and $\hat{\alpha}_{1}^{high}$) are not provided for college-educated parents, as nearly all of them— approximately 98%— have only educated children.
Result 3: Among parents with low financial, social and cultural constraints, the ones without formal education perceive a 6.6% higher utility on high school of education for sons compared to daughters, while for college-educated parents, the utility gap is approximately 2%.
The parameter estimates for parents with low financial constraints indicate those without formal education perceive a 6.6% higher utility on graduating high school for sons compared to daughters, while for those with college education the difference is approximately 2.2%. The cost difference between providing a given level of education to firstborn children compared to second born is very small. For both non-educated and college-educated parents the estimates of the marginal costs different are not statistically different from 0 at 5% significance level. Parents’ perceived average utility at high school level of education is $\approx$ .9% (resp. 0.3%) higher for a second born child compared to firstborn child of the same gender among non-educated parents (resp. college educated parents).
5.2 The Model’s Fit
In this section, I access the fit of the model. Using the estimated parameters $\hat{\theta}$, I solve the household maximization problem, to obtain the optimal distribution $q^{\star}_{h}$ of $q_{T_{h}}$ among children within each household $h$ across a simulated sample of $H$ households. I then compare $q^{\star}$ with the observed data to evaluate the model’s fit. First, I compute the intra-household educational difference between daughters and sons’ educational attainment, as well as between firstborn and second-born children using $q^{\star}$ for the simulated households. The empirical distributions of these within-household differences are compared with the ones from actual educational attainment observed in the data. Figure 11 provides a visual comparison of the model’s distributions with the observed data, showing a clear match. Second, I use $q^{\star}$ to derive targeted moments for estimation, alongside selected non-targeted moments. These analytical outcomes are summarized in Figure 23. It shows that there is no significant difference between the model and the data for all the targeted moments and most of the non-targeted moments as well.
(a) Firstborn vs. second born
(b) Daughter vs. son
Figure 11: Empirical distribution of key moments: Data vs. Model ((For non-educated parents))
5.3 Counterfactual Analysis
Counterfactual 1: Without gender and birth order effects, the distribution of the average educational attainment difference between daughters and sons exhibits first-order stochastic dominance over the distribution in cases with such effects.
In the absence of any gender and birth order effects, differences in education within a household between daughters and sons primarily result from variations in their individual unobserved characteristics/abilities. If children’s innate abilities are assumed to be distributed independently of gender and birth order, the average educational difference between daughters and sons, without considering gender or birth order effects, follows a symmetric distribution centered around 0. However, this distribution shifts towards the negative side in instances when gender and birth order disadvantages are present. In other words, without gender and birth order effects, the distribution of the average education difference between daughters and sons exhibits first-order stochastic dominance over the distribution in cases with such effects (see Figure 12 for non-educated parents). The distribution with no gender and birth order effects can be achieved through the substitution effect of a targeted educational cost reduction policy. Note however that such scenario did not change significantly the overall average within-family disparities in children’s educational attainment among non-educated parents.
(a) Daughter v.s son
(b) Firstborn v.s second born
Counterfactual 2: At the intensive margin, the average educational attainment difference between firstborn daughters and second born sons is equal to 0 in the presence of gender and birth order effects when the firstborn daughter’s ability draw is $\approx 13\%$ (resp. 8%) higher than the ability draw of the second born son for households with non-educated head of household, (resp. households with college educated head of household).
In this counterfactual analysis, my primary goal is to assess how the unobserved source of inequality interact with the observed sources in the educational attainment choice within-family. I do this by quantifying the additional ability needed to counterbalance educational inequality due to gender and birth order effects. First, I compute the extra ability needed by daughters and older siblings to offset the effect of gender and birth order on their educational attainment. In order to do that I solve the household maximization problem in equation 9 with (using $\hat{\theta}$) and without (setting $\theta=0$) gender and birth order effects for a grid of relative ability of children for two-child families with a firstborn daughter and a second-born son, and compute the following quantities:
  1. 1.
    Ability of the firstborn daughter relative to the second born son at which the average difference between daughter’s and son’s education is equal to zero even in the presence of gender and birth order effects.
  2. 2.
    The change in inequality due to gender and birth order effects, by level of relative ability of the firstborn daughter.
Figure 13 presents the results for non-educated and college educated parents with low financial, social and cultural constraints. It suggests two main conclusions. First, for the same ability draws, gender and birth order effects reduce the education attainment of the first born daughter by $\approx 2.2$ years and 1.2 years for non-educated parents and college educated parents respectively. Second, the average education difference between firstborn daughters and second born sons is equal to 0 in the presence of gender and birth order effects when the firstborn daughter’s ability draw is $\approx 13\%$ (resp. 8%) higher than the ability draw of the second born son for households with non-educated head of household, (resp. households with college educated head of household).
(a) Non-educated head of household
(b) College educated head of households
Figure 13: Effect of gender and birth order on inequality ($N_{c}=2$)
5.3.1 Counterfactual 3
This counterfactual is analyzed only among non-educated parents only, and the outcomes are compared to college-educated parents’ observed outcomes.
Barriers to School Entry for all Children: In Theory
The objective of this section is to examine the effectiveness of a counterfactual focusing on removing barriers to school entry, primarily addressing obstacles arising from parental decisions, to ensure that every child is enrolled in the school system. Theoretically this means setting $\nu_{h}$ to 0 for all households. Figure 14 displays the distribution of the difference in education between daughters and sons in panel (a), and between firstborn and second-born children of the same gender in panel (b), across three distinct situations. The elimination of barriers to school entry reduces gender and birth order effects and overall average within-family disparities in children’s educational attainment. In particular, the distribution of the difference between daughters’ and sons’ (resp. firstborn and second born children’s) educational attainments, after removing barriers to school entry for all children, second order stochastically dominates both the distributions with and without gender (and birth order) effects. This means that, compared to the situations there is no gender and birth order effects, removing barriers to school entry leads to more favorable and equitable educational outcomes. The overall distribution shifts in a way that is consistently better, resulting in a notable reduction in average inequality in the sample.
(a) Daughter v.s son
(b) Firstborn v.s second born
Figure 14: Distribution of the difference in children’s education for non-educated parents ($N_{c}=2$) [Observed vs. with no barriers to school entry for all children]
In summary removing barriers to school entry for all children reduces part of the gender and birth order effects on the within-household educational inequality. In addition, overall average within-household inequality is reduced by 50%. However, a significant share ($\approx 40\%$) of the average inequality is attributable to gender effect. In particular, it reduces the gender effect by 60% and birth order effect by 78%. That is consistent with the fact that gender effect is equally present in both the extensive and intensive margin where as the birth order effect is mostly present in the extensive margin.
Barriers to School Entry for all Children: In Practice
In the previous section, I set the financial, social and cultural constraint parameter to low for all households and analyze how within-family disparities in children’s educational attainment is affected. In this section, I analysis how this theoretical situation can be achieved with practical policy tools. Recall that financial, social and cultural constraints are determined by parents’ observable characteristics such as their area of residence (rural vs. urban), their sector of occupation (agricultural vs. non-agricultural), their religion (Christian vs. non-Christian), their level of wealth (low vs. high), and the average educational attainment of their children (a proxy for educational resources available). Some of these characteristics can be changed through a policy intervention. In particular, the financial constraints—wealth level and education resources. Observed data evidence show that children of college-educated parents with two adult children have an average education level of 14.5 years, compared to 9.2 for children of non-educated parents. In addition, average HWI [Household Wealth Index] is 0.03 among non-educated parents compared to 1.06 among college educated parents. In this section, I increase these two quantities for non-educated parents to the same level as college educated parents. The results show a decrease by 10% in the average disparities in children’s educational attainment within-family— not nearly as low as college-educated parents’ level— with 51% of the remaining inequality due to gender effects (see Figure 15). What happened is that the increase in income and educational resources was successful at removing barriers to school entry for children with non-educated parents [we observed a significant increase in the proportion of non-educated children], however, it also created more room for gender bias to create inequality in the intensive margin. And that is why it was not as effective as expected. In summary, some combination of awareness campaigns or laws promoting gender equality in education and an increase in households’ educational resources would be effective.
Figure 15: Effects of increase in income and educational resources
The following Table summarize the effectiveness of these different education policies counterfactual.
Table 7: Summary of Education Policies Counterfactual
Eliminate gender & Reduce average
birth order effects inequality
Remove Gender and birth order disadvantages (1)
Remove barriers to school entry (2)
Increase in income and educational resources (3)
6 Conclusion
In this paper, I examine the interaction between the three empirically known sources of disparities in children’s educational attainment within-households. I propose and estimate a structural model of households’ distribution of education resources among children, allowing for the influence of factors such as gender, birth order and ability of children. The model not only allows me to decompose, for each relative ability draw, the total observed inequality into parts due to gender, birth order, and ability differences; it also gives a platform for analysing how different education policies affect within-household inequalities.
The construction of the model is motivated by contexts similar to the one of Benin; a setting marked by notable disparities in children’s education within-households, coupled with evidence of gender and birth order disadvantages. To ensure tractability, certain aspects of the parental decision-making process regarding education resources distribution are omitted. Notably, the model adopts a static approach, although the education decision of children is inherently dynamic. The primary objective of the paper being to rationalize the observed differences in children’s education, attributing them to gender effect, birth order effect, or variations in innate ability draws; despite its static nature, the model proves relevant, as it effectively incorporates and analyzes the interactions among these three factors. Additionally, the paper attributes any unexplained differences in children’s education, not accounted for by gender and birth order, to differential draws of innate ability. However, it acknowledges the potential influence of other unobserved factors, such as varying favoritism for wives in polygamous households, which could lead to increased parental investment in the education of specific children. In recognizing this, the interpretation of unexplained inequality within-households is acknowledged as an upper bound of the effect of differential ability.
In light of the findings in this paper, we can expect a reduction in the opportunity cost of girls education such as education support in the form of cash transfers, scholarships, and school kits for girls; to reduce within-household inequality in children’s education that is due to gender disadvantage. Additionally, a reduction in the opportunity cost of education for firstborn, such as cash transfers and school kits, to young parents (first-time parents) or scholarships for firstborn children; is expected to reduce within-household inequality in children’s education that is due to birth order disadvantage. However, these two policies need to be combined for an effective reduction in disadvantaged-based inequality. This is due to the possibility of displacement of disadvantage from one group to another. In particular, if the policy only targets firstborn children, the disadvantage against daughters might increase, and vice versa. Finally, a compulsory education policy is the most effective in reducing average inequality in the sample. However, as long as there is budget constraint, as we move toward maximum education for everybody, there will always be a positive within-household inequality.
7 Tables and Figures
Figure 16: Distribution of inequality by number of children.
Figure 17: Within-household standard deviation of children’s education as function of average education of children (Min vs. Max vs. Observed for $N_{c}=2$).
(a) By gender
(b) By birth order
(a) Gender composition of households as function of number of uneducated children
(b) Number of uneducated children as function of gender composition of households
Figure 19: Number of uneducated children by gender composition ($N_{c}=2$)
(a) Estimate
(b) Predicted probability as function of $\bar{q}$
Figure 20: Estimates of $\hat{\beta}$
Figure 22: Plots of the objective function
(a) Average range
(b) Standard deviation
Figure 23: Data vs Model moments (For non-educated parents)
Figure 24: Average Inequality after subsequent shot down of sources of inequality
(a) Mean
(b) Standard Deviation
Figure 26: Mean and standard deviation of adult between 25 and 40 years old
(a) Women
(b) Men
Figure 27: Empirical cdf of the years of education of adults between 25 and 40 years old
Tables
Table 8: Educational inequality on average education
(1) (2) (3) (4) (5) (6)
Full sample Non-educated hh College educated hh
Average education ($\bar{q}$) $1.25^{*}$ $1.04^{*}$ $1.44^{*}$ $1.26^{*}$ $0.87^{*}$ $0.64^{*}$
$[1.24;1.27]$ $[1.02;1.07]$ $[1.42;1.46]$ $[1.23;1.29]$ $[0.81;0.94]$ $[0.54;0.75]$
$\bar{q}^{2}$ $-0.08^{*}$ $-0.06^{*}$ $-0.09^{*}$ $-0.08^{*}$ $-0.05^{*}$ $-0.04^{*}$
$[-0.08;-0.08]$ $[-0.06;-0.06]$ $[-0.09;-0.09]$ $[-0.08;-0.08]$ $[-0.05;-0.04]$ $[-0.04;-0.03]$
R2 $0.69$ $0.71$ $0.73$ $0.74$ $0.61$ $0.62$
Covariates
Num. obs. $32729$ $32729$ $19558$ $19558$ $1438$ $1438$
Null hypothesis value outside the confidence interval. Covariates include parents’ education, area of residence, religion, number of children,
and gender composition of the household.
8 Appendix A: Robustness of Estimates to Missing Siblings
The main sample used for this analysis comprises adult children who were living in the same household as their parents during the census period. This sample represents a specific subgroup within the larger population of adult children. Importantly, the decision for children to leave the parental home is often influenced by factors such as their occupation and educational accomplishments, making it an endogenous process. Moreover, the motives for leaving home frequently differ between daughters, commonly associated with marriage, and sons. Given these dynamics, there’s a potential for bias in our estimates. This would be particularly concerning if, firstly, the children who remained at home are more similar to each other, and secondly, if they significantly differ from those who moved out. The wide range in both educational attainment and gender among children residing in the same household as their parents suggests that the first concern may not be significant.
The second concern could lead to either overestimation— if women who moved out are more educated compared to ones who stayed and men who moved out are less educated compared to the ones who stayed— or underestimation— if women who moved out are less educated and men who moved out are more educated, compared to those who remained at home.
In this section, I delve into the potential bias in estimating the effect of gender disadvantage on within-household inequality. To investigate this, I compare the educational attainment of adult women and men living in the same households as their parents to those who have moved out. The mean comparison between these two groups is presented in Figure 26. This comparison suggests that the difference in average education between men and women is more pronounced in the sub-sample that is not included in my analysis. In addition we observe a clear first order stochastic dominance between the empirical distribution of the education of adult female living in the same households as their parents and those who do not (see Figure 27). Such first order stochastic dominance is not as pronounced among men. As a result, it implies that, if anything, I may be underestimating the effect of gender disadvantage. Consequently, my estimate of gender disadvantage can be interpreted as an estimate of the lower bound of the true parameter.
A similar argument to the one presented in the previous section also applies to the birth order disadvantage parameter. The decision for children to move out is closely linked to their age, with older children being more inclined to leave their parents’ household. Consequently, we may have a selected sample of younger children in some households. In specific cases, children referred to as firstborns in certain households might actually be of a higher birth order. Additionally, more accomplished younger siblings may have already moved out. It’s important to note that both of these situations would potentially bias our estimate of the birth order disadvantage parameter downward. In particular, if we maintain the assumption that firstborn children receive less education than other children, the older firstborn children with less education— who already moved out of the family house— are not included in our analysis. This leads to an underestimation of the birth order effect. In addition if high educated children are more likely to move out— because they have better and stable socio-economic status— we observe uniformly less educated children in our sample. In summary, we are likely to have in our sample, less educated children. On one hand, if the age effect dominates the education effect, we will have less firstborn children in our sample, which biases our estimate downward. On the other hand, if the education effect dominates the age effect, we have less second born children in our sample, which also biases our estimate downward.
Appendix B
8.1 Appendix B1: Summary Statistics
Table 9: Summary Statistics for all the individuals between 25 and 40 years old residing in the same household as their parents
Mean Standard Deviation
All 5.11 5.94
Rural 3.42 5.18
Urban 7.05 6.14
Farmer 2.91 4.83
Education Non-farmer 6.75 6.15
Non-Christian 3.43 5.23
Christian 7.13 6.1
Non-educated parents 3.45 5.16
College-educated parents 13.8 4.42
No education 0.48
Female 0.38
Rural 0.53
Farmer 0.43
Christian 0.45
Non-educated parents 0.72
College-educated parents 0.03
HWI 0.03 0.98
Number of children 5.35 4.1
Number of children btw 25 and 40 2.1 1.51
Average educ in hh is 0 0.4
hh with only one child btw 25 and 40 0.45
Number of observation 253,734
8.2 Appendix B2: Model for households with Number of Children equal 3
For $N_{c}=3$, let $ne\in\{0,1,2\}$ be number of children with no formal and education, and $\nu_{h}^{L}=1\{ne_{h}=0\}$, $\nu_{h}^{M}=1\{ne_{h}=1\}$, and $\nu_{h}^{H}=1\{ne_{h}=2\}$
$U(q_{h},\theta)=\nu_{h}^{L}\Big{[}\sum_{i=1}^{N_{c}}a_{i}.(q_{i})^{\delta_{i,h}^{L}}-\alpha_{i}^{L}q_{i}\Big{]}+\nu_{h}^{M}\Big{\{}\sum_{i=1}^{N_{c}}\Big{[}e_{i}^{M}.\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{M}}-\alpha_{i}^{M}q_{i}\big{)}\Big{]}\Big{\}}+\nu_{h}^{H}\Big{\{}\sum_{i=1}^{N_{c}}\Big{[}e_{i}^{H}.\Big{[}\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{H}}-\alpha_{i}^{H}q_{i}\big{)}\Big{]}\Big{\}}$ (12)
where,
$\delta_{i,h}^{type}=\gamma-\theta_{1}^{type}Female_{i}\frac{1}{N_{c}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j}),\text{ with }type\in\{low(L),medium(M),high(H)\}$
$e_{i}^{M}=1\{\exists j:\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{M}}-\alpha_{i}^{M}q_{i}\big{)}>\big{(}a_{j}.(q_{i})^{\delta_{j,h}^{M}}-\alpha_{j}^{M}q_{j}\big{)}\},$
$e_{i}^{H}=1\{\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{H}}-\alpha_{i}^{H}q_{i}\big{)}>\big{(}a_{j}.(q_{i})^{\delta_{j,h}^{H}}-\alpha_{j}^{H}q_{j}\big{)},\forall j\neq i\},$
The vector of parameters of interest is
$\theta=\Big{(}\theta_{1}^{L},\theta_{1}^{M},\theta_{1}^{H},\alpha_{(1)}^{L},\alpha_{(2)}^{L},\alpha_{(1)}^{M},\alpha_{(2)}^{M},\alpha_{(1)}^{H},\alpha_{(2)}^{H}\Big{)}$
$\alpha_{(3)}^{L}$, $\alpha_{(3)}^{M}$, $\alpha_{(3)}^{H}$ are normalized to 0.The dimension of $\theta$ is $1\times 9$.
For each household $h$, [Parents’ educ and number of children hold fixed], let
  • $\nu_{hne}^{\star}$ be the utility [due to unobserved financial, social & cultural constraints] from having $ne$ children with no formal education.
  • $ne\in\{0,1,\dots,N_{c}-1\}$ be the observed discrete outcome of number of children with no formal educ.
$\nu_{hne}^{\star}=X_{h}^{\prime}\beta_{ne}+\varepsilon_{hne}$
[$X_{h}$ includes a cst, rural, farmer, HWI, religion, gender composition, and average education of children.]
The choice of children with no formal education is
$ne\text{ if }\nu_{hne}^{\star}>\nu_{hne^{\prime}}^{\star}\forall ne^{\prime}\neq ne$
$ne_{h}=argmax_{ne}X_{h}^{\prime}\hat{\beta}_{ne}+\varepsilon_{hne}^{Sim}\text{, with }\varepsilon_{hne}^{Sim}\sim\mathcal{N}(0,\hat{\Sigma})$
Table 10: Estimates of $\hat{\theta}$, ($N_{c}=3$)
$\hat{\theta}_{1}^{L}$ $\hat{\alpha}_{1}^{L}-\hat{\alpha}_{2}^{L}$ $\hat{\alpha}_{2}^{L}-\hat{\alpha}_{3}^{L}$ $\hat{\theta}_{1}^{M}$ $\hat{\alpha}_{1}^{M}-\hat{\alpha}_{2}^{M}$ $\hat{\alpha}_{2}^{M}-\hat{\alpha}_{3}^{M}$ $\hat{\theta}_{1}^{H}$ $\hat{\alpha}_{1}^{H}-\hat{\alpha}_{2}^{H}$ $\hat{\alpha}_{2}^{H}-\hat{\alpha}_{3}^{H}$
Estimates 0.033∗∗ 0.0026∗∗ 0.0010 0.0761∗∗ 0.0094 0.0001∗∗ 0.1596∗∗ 0.0120∗∗ 0.0105∗∗
Standard errors 0.0052 0.0006 0.0006 0.0058 0.0016 0.0019 0.0073 0.0019 0.0015
Number of observations 3644
**∗ ∗ significant at 5% level of significance, * significant at 10% level of significance.
8.3 Appendix B3: Generalized Households’ utility function for any Number of Children
Let $type\in\{0,1,\dots N_{c}-1\}$, where
  • 00 corresponds to the least constrained household,
  • and $N_{c}-1$ corresponds to the most constrained household.
$U(q_{h},\theta)=\sum_{c=2}^{N}\Big{\{}1\{N_{c_{h}}=c\}.U^{c}(q_{h},\theta)\Big{\}}\text{ where,}$, (13)
  • $\delta_{i,h}^{type}=\gamma-\theta_{1}^{type}Female_{i}\frac{1}{N_{c_{h}}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j})$,
  • $q_{h}=(q_{1,h},\dots,q_{N_{c},h})$
9 Appendix C: Additional Figures